OpenStudy (anonymous):

Algebra 2 Find the quotient. 2x-3/x/7/x^2 A. 7/x(2x-3) B. 7x/2x-3 C. 2x-3/7x D. x(2x-3)/7

Nnesha (nnesha):

\[\huge\rm \frac{ \frac{ 2x-3 }{ x } }{ \frac{ 7 }{ x^2 } }\] like this ?

OpenStudy (anonymous):

so the answers is C @Nnesha

Nnesha (nnesha):

how did you get it ?

OpenStudy (anonymous):

I just saw what u typed me and i figure out that is was exacly urs :) @Nnesha so I'm right?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\huge\rm \frac{ \frac{ 2x-3 }{ x } }{ \frac{ 7 }{ x^2 } }\] like this ? \(\color{blue}{\text{End of Quote}}\) i just asked a question is this ur question ?

OpenStudy (anonymous):

yes is like that the question I'm asking @Nnesha

Nnesha (nnesha):

okay change division to multiplication multiply the top with the `reciprocal` of the bottom fraction

OpenStudy (anonymous):

so i see that the answer is C @Nnesha right?

Nnesha (nnesha):

please show ur work so i can check !

OpenStudy (anonymous):

I did this 2x-3/7x cuz I cancel out the x^2

OpenStudy (anonymous):

I really dont know how to slove it :( @Nnesha

Nnesha (nnesha):

here is an example \[\huge\rm \frac{ \color{Red}{\frac{ a }{ b }}}{ \frac{ c }{ d} }= \color{ReD}{\frac{a}{b}} \times \frac{d}{c}\]

Nnesha (nnesha):

what is the reciprocal of the bottom fraction ??

OpenStudy (anonymous):

x and 7

Nnesha (nnesha):

what's reciprocal of 7/x^2 ??

Nnesha (nnesha):

reciprocal of 3/2 is 2/3

OpenStudy (anonymous):

7/x^2

Nnesha (nnesha):

that's the original fraction what's reciprocal of that ?

OpenStudy (anonymous):

i mean x^2/7 @Nnesha

Nnesha (nnesha):

yes right!

Nnesha (nnesha):

so multiply \[\huge\rm \frac{ 2x-3 }{ x } \times \frac{ x^2 }{ 7}\] now multiply the numerator by numerator and denominator by denominator

OpenStudy (anonymous):

I got C

Nnesha (nnesha):

gtg sorry no it's not c

OpenStudy (anonymous):

So D?